Given; 2x+6(x-3)=150
Prove: x=21

Answer by fractalier(6550)

From 2x+6(x-3)=150
we plug 21 in for x and get
2(21) + 6(21-3) = 150
42 + 6(18) = 150
42 + 108 = 150
150=150 and it checks